The length of rectangle is 8 cm more than twice its width. If the area is 960 cm^2. Find its length and width?
Answer by graphmatics(170)

Let x be the width of the rectangle. We are given that 2x+8 is the length of the rectangle. As the area is 960 cm^2 we have that:
x(2x+8)=960
2x^2+8x=960
x^2+4x-480=0
(x-20)(x+24)=0
x=20, x=-24
So we must select x=20 as the width of the rectangle with 2x+8=48 as the rectangle length.